# Galled Perfect Transfer Networks

### v1

@@ 5 Galled PTNs on functional characters: a case study
We now turn to a real-data case analysis of galled completion in order to gain better insight on its potential for the prediction horizontal gene transfers. Its applicability is not fully understood, mainly due to the lack of benchmarks for the interactions between transfers and characters. However, a few studies have now started exploring this territory, and here we take inspiration from [machine], which implemented machine learning approaches to infer HGTs from functional characters using the Kyoto Encyclopedia of Genes and Genomes (KEGG) Orthology database [KEGG] (see below). The authors used curated HGTs between bacteria and the list of functions of each taxa as features, allowing them to predict the presence or absence of HGT between any two bacterial species. They inferred 147,889 transfers between 6566 bacterial genomes, although some of these transfers could all be manifestations of one ancestral transfer in the phylogeny (but because these predictions are made in a pairwise manner, the time of these transfers cannot be inferred by the approach).
We also used functional characters from the KEGG database, which associates molecular functions of genes and proteins to orthologous groups. Table 1 lists the functions that we used, along with their hierarchical classification. The main type of entries in the database are KEGG Orthologs (KO), which are groups of genes sharing common functions (members of the same group are viewed as functional orthologs). Most groups are formed based on experimental evidence, with some generalizations to other organisms based on sequence similarity.
We took a subset of 45 species from the bacterial species used in [machine], which consists of species that were predicted to be involved in interphylum transfers. We obtained the corresponding species tree from NCBI Taxonomy Browser [schoch2020ncbi], noting that it is not completely resolved and is therefore non-binary. The whole annotated genomes of these species are contained in KEGG. As characters, we chose 23 KOs based on the features that shown as most important for the Graph Convolutional Network model presented in [machine]. These features include metabolism-related, information processing and antibiotic resistance KOs among other functions. The whole datasets and scripts used for this part are contained in the following repository: https://github.com/AliLopSan/ptns.
@@
Figure 9: Distribution of the number of first-appearances per character. We distinguish those first-appearance nodes that are leaves and those that are inner nodes. Note that in general characters have more than two first-appearances, thus if there exists a PTN that can explain them, this is not necessarily galled. We note however that there exists a set of characters that have at most two first appearances:$\{K13628,K19115,K00561,K19310\}$ which are those that fall below the green dashed line.
@@
Figure 11: (a) The species tree with 45 species with annotated leaves showing colored annotations by clade according to the presence of one of the characters: $\{K13628,K19115,K00561,K19310\}$. Note that between annotations of the same color there should exist a transfer edge that could join the parts of the tree where the characters are present. Leaf names correspond to the species id present in the KEGG orthology database. Notice that this network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node. However as shown in (b), by resolving one of the children of the root node, as the black dot, this network becomes galled. Tree annotations where visualized using iTOL tool  [itol].
@@
Using the Fitch labeling algorithm described in [ptn], we first calculated the number of first-appearance nodes of each character as well as their position in the species tree. As established above, this serves as a first test for galled-completability, as characters with three of more such nodes are sufficient to discard this property. Figure 9 shows the distribution of the number of first-appearance nodes for each character. It is immediately clear that the “at most two first-appearances” condition is far from being satisfied, although it is insightful to understand why. For several characters, the majority of first-appearance nodes are leaves. An example of this phenomenon is shown in Figure 10. Notably, there is a clade in the top-right part of the tree in which every leaf has the character except one, resulting in three first-appearances in that clade. This prevents the common ancestor of the clade from being labeled with that character. It is plausible that one character was simply lost in one species, without invoking any transfer, as shown on the right. This strongly indicates that the “transfer-only” model is too rigid, and in future work we plan to work on approaches to identify situations in which losses are more likely than transfers.
@@
Going back to Figure 9, we note that there does exist characters that posses at most two first-appearances. We checked whether this subset of four characters could at least be galled-completable, but it turns out that this is not even the case. This is shown in Figure 11 on the left. As explained in the caption, if we built the redundancy-free network for these characters and applied Lemma 3, we would add a transfer edge linking the branches above the roots of the blue character, and above the roots of the green character. These are non-redundant edges that create two cycles that intersect at the root. But one could hypothesize that the unresolved root is responsible for this: Figure 11 on the right partially resolves the tree by expanding the root, creating a branch in a way that the cycle of the blue characters now avoids the root. One can now check that the resulting redundancy-free network is galled, and that this set of characters is galled-completable.

### v2

@@ 5 Galled PTNs on functional characters: a case study
We now turn to a real-data case analysis of galled-completion in order to gain better insight on its potential for the prediction horizontal gene transfers, as there is a general lack of benchmarks for the interactions between transfers and characters. However, a few studies have now started exploring this territory, and here we take inspiration from [machine], which implemented machine learning approaches to infer HGTs from functional characters using the Kyoto Encyclopedia of Genes and Genomes (KEGG) Orthology database [KEGG] (see below). The authors of  [machine] used curated HGTs between bacteria, with the list of functions of each taxa as features, allowing them to predict the presence or absence of HGT between any two bacterial species. They therefore solve a binary classification task, making one prediction per species pair. They inferred 147,889 transfers between 6566 bacterial genomes, although some of these transfers could all be manifestations of one ancestral transfer in the phylogeny (but because these predictions are made in a pairwise manner, the time of these transfers cannot be inferred by the approach).
In our case study, we also used functional characters from the KEGG database, which associates molecular functions of genes and proteins to orthologous groups. Table 1 lists the functions that we used, along with their hierarchical classification. The main type of entries in the database are KEGG Orthologs (KO), which are groups of genes sharing common functions (members of the same group are viewed as functional orthologs). Most groups are formed based on experimental evidence, with some generalizations to other organisms based on sequence similarity.
We took a subset of 45 species from the bacterial species used in [machine], which consists of species that were predicted to be involved in interphylum transfers. We obtained the corresponding species tree from NCBI Taxonomy Browser [schoch2020ncbi], noting that it is not completely resolved and is therefore non-binary. Observe that this is not a problem for our tree-completion procedure, as only the inserted transfer nodes are assumed to be binary in our model — not the nodes of the input tree. The whole annotated genomes of these species are contained in KEGG, and so as characters, we chose 23 KOs based on the features that were shown as most important for the Graph Convolutional Network model presented in [machine]. These features include metabolism-related, information processing and antibiotic resistance KOs among other functions. The whole datasets and scripts used for this part are contained in the following repository: https://github.com/AliLopSan/ptns.
@@
Figure 10: Distribution of the number of first-appearances per character. We distinguish those first-appearance nodes that are leaves and those that are inner nodes. Most characters have more than two first-appearances, thus if there exists a PTN that can explain them, it will not be galled. We note however that there exists a set of characters that have at most two first appearances:$\{K13628,K19115,K00561,K19310\}$ which are those that fall below the green dashed line.
@@
Figure 12: (a) The species tree with 45 species with annotated leaves showing colored annotations by clade according to the presence of one of the characters: $\{K13628,K19115,K00561,K19310\}$. Note that between annotations of the same color there should exist a transfer edge that could join the parts of the tree where the characters are present. Leaf names correspond to the species id present in the KEGG orthology database. Notice that the redundancy-free network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node. However as shown in (b), by resolving one of the children of the root node, as the black dot, this network becomes galled. Tree annotations where visualized using iTOL tool  [itol].
Losses can obstruct galled-completability.
@@
Using the Fitch labeling algorithm described in [ptn], we first calculated the number of first-appearance nodes of each character as well as their position in the species tree. As established in Lemma 3, this serves as a first test for galled-completability, as finding three or more such nodes for a character is sufficient to discard the instance. Figure 10 shows the distribution of the number of first-appearance nodes for each character. It is immediately clear that the “at most two first-appearances” condition is far from being satisfied, although it is insightful to understand why. For several characters, the majority of first-appearance nodes are leaves. An example of this phenomenon is shown in Figure 11. Notably, there is a clade in the top-right part of the tree in which every leaf has the character except one, resulting in three first-appearances in that clade. This prevents the common ancestor of the clade from being labeled with that character. It is plausible that the character was transferred to the common ancestor of the clade, and then simply lost in one species, as shown on the right. This strongly indicates that the “transfer-only” model is too rigid, and in future work we plan to work on approaches to identify situations in which losses are more likely than transfers.
Unresolved species trees can obstruct galled-completability.
@@
Going back to Figure 10, we note that there do exist characters that possess at most two first-appearances. We checked whether this subset of four characters could at least be galled-completable, but it turns out that this is not even the case. This is shown in Figure 12 on the left. As explained in the caption, if we built the redundancy-free network for these characters and applied Lemma 3, we would add a transfer edge linking the branches above the roots of the blue character, and above the roots of the green character. These are non-redundant edges that create two cycles that intersect at the root. But one could hypothesize that the unresolved root is responsible for this: Figure 12 on the right partially resolves the tree by expanding the root, creating a branch in a way that the cycle of the blue characters now avoids the root. One can now check that the resulting redundancy-free network is galled, and that this set of characters is galled-completable.